# KG-Hopper: Empowering Compact Open LLMs with Knowledge Graph Reasoning via Reinforcement Learning

### 22 Mar 2026 v1

@@ I Introduction
Large Language Models (LLMs) have achieved remarkable success across diverse domains but still exhibit notable shortcomings, such as hallucinations and factual inaccuracies, particularly in knowledge-intensive tasks [wang2022knowledge]. This is largely due to the implicit storage of knowledge within their model weights, making knowledge updates cumbersome and resource-intensive through fine-tuning [wang2025domain]. To address this limitation, Retrieval-Augmented Generation (RAG) has emerged as an effective strategy, enabling LLMs to dynamically access external knowledge during inference [wang2025plugging]. Among various external resources, Knowledge Graphs (KGs) stand out as structured and reliable knowledge bases, offering explicit, interpretable, and easily updateable knowledge beneficial in critical applications such as medicine and finance [wang2024knowledge].

### 27 Mar 2026 v2

@@ I Introduction
Large Language Models (LLMs) have achieved remarkable success across diverse domains but still exhibit notable shortcomings, such as hallucinations and factual inaccuracies, particularly in knowledge-intensive tasks. This is largely due to the implicit storage of knowledge within their model weights, making knowledge updates cumbersome and resource-intensive through fine-tuning. To address this limitation, Retrieval-Augmented Generation (RAG) has emerged as an effective strategy, enabling LLMs to dynamically access external knowledge during inference [wang2025plugging]. Among various external resources, Knowledge Graphs (KGs) stand out as structured and reliable knowledge bases, offering explicit, interpretable, and easily updateable knowledge beneficial in critical applications such as medicine and finance.